OpenStudy (anonymous):

Hi I need help on a few problems. I have been trying to do them but I can't seem to get the correct answer. 1. A population P is initially 2000. Find an exponential model (growth or decay) for the population after t years if the population P is increases by 400% every 3 years. (Round your terms to two decimal places.) P(t)= ? 2. A population P is initially 2000. Find an exponential model (growth or decay) for the population after t years if the population P is multiplied by 0.85 every month. (Round your terms to four decimal places.) P(t)= ???

jimthompson5910 (jim_thompson5910):

For #1, you're using the formula A = P(1+r)^(t/k) where A = final population at time t P = initial population r = growth rate (in decimal form) t = time in years k = frequency when the population increases So in this case A = unknown P = 2000 r = 4 (400% = 4) t = unknown k = 3 (the frequency is every 3 years)

jimthompson5910 (jim_thompson5910):

does that help?